# Spectral Structure of the Mixed Hessian of the Dispersionless Toda $\tau$-Function

### 24 Mar 2026 v1

@@ 1.2. Organization of the paper
Section 2 analyzes the inverse conformal map, derives the Raney-number formulas, and identifies the square-root singularity at the analytic threshold. Section 3 introduces the scale-invariant Toda Hessian, proves the blockwise Gram factorization, and records the scalar Gram quantities that will later be continued across the threshold. Section 4 develops the weighted spectral theory in the subcritical regime $0<\zeta<\zeta_{c}$ and proves Theorem A. Section 5 studies the scalar continuation problem beyond $\zeta_{c}$, including the hypergeometric representation, the resonant expansion, the Cauchy–Stieltjes formula, and the Jacobi realization. Finally, Section 6 treats the geometric threshold $\zeta_{\mathrm{univ}}$, proves the strict separation (1.7), and shows that the continued scalar data remain finite at the univalence boundary. The appendices contain the uniform Raney asymptotics, the entrywise rank-one extraction, the hypergeometric continuation argument, and the computation of the branch-point coefficient.
@@ 2. Inversion of the conformal map
This section analyzes the inverse conformal map and extracts the information needed for the spectral analysis of the Hessian. It provides three inputs for later use: the functional equation for the inverse branch (Proposition 2.1), the explicit Taylor coefficients in terms of Raney numbers (Proposition 2.4), and the location and square-root nature of the dominant singularity (Lemma 2.2). The last point generates the universal $m^{-3/2}$ coefficient decay responsible for the logarithmic spectral divergence established in Section 4.
@@ 2.1. The generating function and its functional equation
We encode the inverse branch $w(z)\sim z/r$ as $z\to\infty$ by factoring out the leading linear growth:
@@
Let $f(w)=rw+aw^{1-s}$ with $\zeta=a/r$, and let $w(z)$ be the inverse branch normalized by $w(z)\sim z/r$ at infinity. Define $U(x;\zeta)$ by (2.2). Then $U$ satisfies the algebraic functional equation
@@ 2.2. Square-root criticality
We now determine the local behavior of $U(x;\zeta)$ at its dominant singularity. This is the key analytic input for the spectral analysis: the square-root nature of the singularity produces the universal $m^{-3/2}$ coefficient asymptotics that drive the logarithmic divergence of the Hessian.
@@
The function $U(x;\zeta)$ defined by $U=1+\zeta x^{s}U^{s}$ is a power series in the combination $\zeta x^{s}$, with radius of convergence $\zeta_{c}=(s-1)^{s-1}/s^{s}$ in that variable. Equivalently, for fixed $\zeta>0$ its radius of convergence in $x$ is $(\zeta_{c}/\zeta)^{1/s}$. Its only singularity on $|\zeta x^{s}|=\zeta_{c}$ is at $\zeta x^{s}=\zeta_{c}$. At this point the function has a square-root branch point with the local expansion
@@ 2.3. Raney numbers and coefficient formulas
The Taylor coefficients of powers of $U$ admit a classical closed form. We define these coefficients directly and then identify them with the Raney numbers.
For any $p\in\mathbb{Z}$, write
@@
For $p\geq 1$, these are the classical Raney numbers; the case $p=1$ gives the Fuss–Catalan numbers [20, 8].
@@
We now record the expansions needed for the Hessian kernel in Section 3. The kernel involves the logarithm $\log(1-(w\bar{w}^{\prime})^{-1})$, which we will expand in powers of $w^{-1}$; hence we need the following.
@@
with an explicit amplitude $A_{s,p}>0$ depending on $s$ and $p$. The exponent 3/2-3/2 is universal for simple algebraic singularities of square-root type. This is the only asymptotic information about the inverse map that enters the Hessian estimates in Section 4; the detailed derivation is given in Appendix A.
@@ 3. The Hessian of the $\tau$-function
This section constructs the main object of the paper: the mixed Hessian of $\log\tau$ with respect to the dispersionless Toda times. We work throughout in the scale-free variables introduced in Section 2, which removes the trivial dependence on the conformal radius and yields a Hessian that depends only on the shape parameter $\zeta=a/r$.
The section is organized as follows. We first recall the kernel representation due to Wiegmann–Zabrodin and use it to define the scale-invariant Hessian (Subsection 3.1). We then expand the kernel using the Raney machinery from Section 2, obtaining an exact Gram representation as a sum of rank-one operators (Subsection 3.3). Finally, we analyze the Gram weights and identify the analytic threshold $\zeta_{c}$ at which the naive unweighted Gram realization breaks down (Subsection 3.4).
@@ 3.1. Kernel representation and scale-invariant formulation
The dispersionless Toda hierarchy associates to a simply connected domain $D\subset\mathbb{C}$ a $\tau$-function whose logarithm $\mathcal{F}=\log\tau$ serves as a generating function for the harmonic moments. In the present paper, however, we use this framework only for the polynomial family (1.3). The general discussion from Section 1 is included only to place the problem in its geometric and integrable context.
For the maps (1.3) with $0<\zeta<\zeta_{\mathrm{univ}}$, the boundary is real-analytic and the Wiegmann–Zabrodin kernel identity applies. The mixed second derivatives of $\mathcal{F}$ with respect to the Toda times $(t_{m},\bar{t}_{n})$ are then expressed via the inverse conformal map $w(z)$ as follows [13]:
@@
The scale-invariant Hessian $H=(H_{mn})_{m,n\geq 1}$ is naturally a positive semidefinite quadratic form on finitely supported sequences. Spectral statements require a compatible Hilbert realization, and the standard $\ell^{2}(\mathbb{N})$ topology is not stable at analytic criticality. The appropriate framework is the weighted Gram realization introduced later in Definition 4.1; all spectral statements in Sections 45 refer to these renormalized operators.
@@ 3.3. Gram representation
Then the scale-invariant Hessian admits the exact representation
@@
where $\bm{v}\otimes\bm{v}^{*}$ denotes the rank-one operator $(\bm{v}\otimes\bm{v}^{*})_{mn}=v_{m}\overline{v_{n}}$.
@@
The Gram form (3.9) makes positivity of $H$ manifest: it is a sum of positive semidefinite rank-one operators. More importantly, it separates two sources of complexity: (i) The individual contribution of each logarithmic mode $p$, encoded in the vector $\bm{v}^{(p)}$, and (ii) the collective mixing of modes within each symmetry sector, which determines the eigenvalue distribution.
The spectral behavior of $H$ is controlled by the interplay between these two effects. As we show next, the critical phenomenon at $\zeta=\zeta_{c}$ arises from the borderline summability of the Gram data, and not from any individual term becoming singular.
@@ 3.4. Gram weights and the analytic threshold
For each fixed $p\geq 1$, one has $\sigma_{p}(\zeta)<\infty$ if and only if $\zeta<\zeta_{c}$. In particular, the analytic threshold for the scalar Gram weights is exactly $\zeta_{c}$.
@@ 4. Subcritical phase ($0<\zeta<\zeta_{c}$)
This section develops the operator-theoretic picture in the analytic subcritical regime. Proposition 3.5 gives a canonical Gram factorization of each symmetry block of the Hessian. Although the unrenormalized Gram weights $\sigma_{p}(\zeta)$ diverge as $\zeta\uparrow\zeta_{c}$ (Proposition 3.8), one can renormalize the column modes by an explicit diagonal weight. The resulting compact positive operators exhibit a single logarithmically diverging eigenvalue, while the remaining soft spectrum stays bounded and admits a well-defined limiting description after compression.
4.1. Operator setting, block factorization, and weighted renormalization
@@
Let $\mathcal{W}:\mathcal{H}_{\beta}\to\ell^{2}(\mathbb{N}_{0})$ be the unitary map
@@
The renormalized synthesis operator is
@@
initially defined on finitely supported sequences and extended by continuity. Its associated weighted Gram and Hessian operators are
@@
The coefficient matrix of the Hessian block is the intrinsic object. The weights (4.2) do not modify these coefficients and do not remove the critical instability; they only realize the same coefficient data in a fixed, $\zeta$-independent Hilbert scale in which the singular part becomes visible as a compact-plus-rank-one operator.
@@
It is thus enough to bound $\sigma_{p}(\zeta)$ uniformly in $p$.
Using the closed form for $R_{s,p}(m)$ from Proposition 2.4, the Gram weights may be written as
@@
for some finite constant $C_{s}^{\prime}(\zeta)$ depending on $\zeta$, but not on $p$.
Finally, using $w_{j}^{2}=p_{j}^{3+2\beta}M^{2p_{j}}$ and (4.5),
@@ 4.2. Rank-one logarithmic spike at $\zeta_{c}$
The detailed coefficient analysis is contained in Appendix B; here we only record the reduction that produces (4.9).
@@
In particular, $R^{(q)}(\zeta)$ is compact for each $\zeta$, and the convergence also holds in operator norm. We denote this operator-norm limit by $\widetilde{C}_{*}^{(q)}$; thus the $R_{*}^{(q)}$ of Appendix B and the $\widetilde{C}_{*}^{(q)}$ used in the main text are the same operator.
@@
because $R^{(q)}(\zeta)\to\widetilde{C}_{*}^{(q)}$ in Hilbert–Schmidt norm and $L(\zeta)(K_{\eta}-K_{1})\to 0$ in operator norm.
Since $K_{1}=\widetilde{\bm{d}}^{(q)}\otimes\widetilde{\bm{d}}^{(q)*}$, this is exactly (4.9). ∎
@@ 4.3. Spectral asymptotics
We now convert the rank-one decomposition of Theorem 4.6 into spectral information. The conclusion is that the singular part produces one stiff eigenvalue in each symmetry block, while the remaining spectrum stays bounded and converges to the compressed limit operator.
@@
Theorem 4.7 has a clear numerical signature: after weighted renormalization, exactly one eigenvalue in each symmetry block grows on the logarithmic scale $L(\zeta)$, while the remaining eigenvalues stay bounded. Figure 1 illustrates this separation for the block operator $\widetilde{G}^{(q)}(\zeta)$ in the sector $q=1$ for $s=3$ and $s=5$. The top row shows that the leading branch $\mu_{1}^{(q)}(\zeta)$ is asymptotically affine in $L(\zeta)$, whereas the bottom row confirms that after division by $L(\zeta)$ only the stiff branch survives and all soft branches tend to zero.
Figure 1. Logarithmic spectral asymptotics of the weighted Gram block $\widetilde{G}^{(q)}(\zeta)$ for $s=3,5$ in sector $q=1$ ($\beta=1$, $N=30$). Top row: the leading eigenvalue $\mu_{1}^{(q)}(\zeta)$ plotted against $L(\zeta)$, showing the asymptotically affine law $\mu_{1}^{(q)}(\zeta)=\Gamma^{(q)}\,L(\zeta)+O(1)$; the dashed line is a tail linear fit. Bottom row: the normalized eigenvalues $\mu_{k}^{(q)}(\zeta)/L(\zeta)$. The ratio $\mu_{1}^{(q)}(\zeta)/L(\zeta)$ approaches a positive constant, whereas $\mu_{k}^{(q)}(\zeta)/L(\zeta)\to 0$ for $k\geq 2$.
Proof.
@@
The remaining spectral data are governed by the compression of the limiting remainder to the orthogonal complement of the spike direction.
@@
Proposition 4.8 identifies the limiting bounded spectrum that remains after the stiff direction is removed. This bounded part is naturally seen at the level of the compressed operator, not in the full diverging spectrum. Figure 2 illustrates this soft regime for $s=3$ and $s=5$: the upper panels show the convergence of the first soft branches as $\zeta\uparrow\zeta_{c}$, while the lower panels display near-critical finite-$\zeta$ soft spectral profiles across the symmetry sectors. Together with Figure 1, this makes the stiff/soft decomposition visually explicit.
Figure 2. Soft spectrum after removal of the stiff direction for $s=3,5$ ($\beta=1$, $N=40$). Top row: the soft branches $\mu_{2}^{(q)}(\zeta),\dots,\mu_{6}^{(q)}(\zeta)$ in sector $q=1$, plotted against $1/L(\zeta)$. Their flattening as $1/L(\zeta)\to 0$ shows that the soft branches remain bounded as $\zeta\uparrow\zeta_{c}$. Bottom row: finite-$\zeta$ snapshots of the soft branches for all sectors $q=1,\dots,s$ at $\zeta/\zeta_{c}=0.9999$.
@@
5. Continuation of scalar Gram data beyond the analytic threshold
5.1. Intrinsic Hessian, weighted realization, and continued scalar data

### 16 Apr 2026 v2

@@ 1.2. Organization of the paper
Section 2 analyzes the inverse conformal map, derives the Raney-number formulas, and identifies the square-root singularity at the analytic threshold. Section 3 introduces the scale-invariant Toda Hessian, proves the blockwise Gram factorization, and presents the scalar Gram quantities that will later be continued across the threshold. Section 4 develops the weighted spectral theory in the subcritical regime $0<\zeta<\zeta_{c}$ and proves Theorem A. Section 5 studies the scalar continuation problem beyond $\zeta_{c}$, including the hypergeometric representation, the resonant expansion, the Cauchy–Stieltjes formula, and the Jacobi realization. Finally, Section 6 treats the geometric threshold $\zeta_{\mathrm{univ}}$, proves the strict separation (1.7), and shows that the continued scalar quantities remain finite at the univalence boundary. The appendices contain the uniform Raney asymptotics, the entrywise rank-one extraction, the hypergeometric continuation argument, and the computation of the branch-point coefficient.
@@ 2. Inversion of the conformal map
We begin with the inverse conformal map, whose singularity structure drives the later spectral analysis. The main results of this section are the functional equation for the inverse map (Proposition 2.1), the explicit Taylor coefficients in terms of Raney numbers (Proposition 2.4), and the identification of the dominant singularity together with its square-root character (Lemma 2.2). The last point produces the universal $m^{-3/2}$ coefficient decay responsible for the logarithmic spectral divergence established in Section 4.
@@ 2.1. The generating function and its functional equation
We represent the inverse map $w(z)\sim z/r$ as $z\to\infty$ by factoring out the leading linear growth:
@@
Let $f(w)=rw+aw^{1-s}$ with $\zeta=a/r$, and let $w(z)$ be the inverse map normalized by $w(z)\sim z/r$ at infinity. Define $U(x;\zeta)$ by (2.2). Then $U$ satisfies the algebraic functional equation
@@ 2.2. Square-root criticality
We now determine the local behavior of $U(x;\zeta)$ at its dominant singularity. This is the central analytic ingredient for the spectral analysis: the square-root nature of the singularity produces the universal $m^{-3/2}$ coefficient asymptotics that drive the logarithmic divergence of the Hessian.
@@
The function $U(x;\zeta)$ defined by $U=1+\zeta x^{s}U^{s}$ is a power series in the combination $\zeta x^{s}$, with radius of convergence $\zeta_{c}=(s-1)^{s-1}/s^{s}$ in that variable. Equivalently, for fixed $\zeta>0$ its radius of convergence in $x$ is $(\zeta_{c}/\zeta)^{1/s}$. Its only singularity in the variable $t:=\zeta x^{s}$ on $|t|=\zeta_{c}$ is at $t=\zeta_{c}$. Equivalently, for fixed $\zeta>0$ the corresponding singularities in the $x$-plane form the $s$-point orbit $\zeta x^{s}=\zeta_{c}$. At this point the function has a square-root branch point with the local expansion
@@ 2.3. Raney numbers and coefficient formulas
The Taylor coefficients of powers of $U$ admit a classical closed form. We define these coefficients directly and then identify them with the Raney numbers. For any $p\in\mathbb{Z}$, write
@@
For $p\geq 1$, these are the classical Raney numbers. In particular, the case $p=1$ gives the Fuss–Catalan numbers [19, 7].
@@
We now present the expansions needed for the Hessian kernel in Section 3. The kernel involves the logarithm $\log(1-(w\bar{w}^{\prime})^{-1})$, which we will expand in powers of $w^{-1}$. Hence we need the following.
@@
with an explicit amplitude $A_{s,p}>0$ depending on $s$ and $p$. The exponent 3/2-3/2 is universal for simple algebraic singularities of square-root type. This is the only asymptotic information about the inverse map that enters the Hessian estimates in Section 4. The detailed derivation is given in Appendix A.
@@ 3. The Hessian of the $\tau$-function
We now introduce the mixed Hessian of $\log\tau$ with respect to the dispersionless Toda times. Throughout, we use the scale-free variables from Section 2, which remove the trivial dependence on the conformal radius and leave a Hessian depending only on the shape parameter $\zeta=a/r$. We first recall the Wiegmann–Zabrodin kernel formula and define the scale-invariant Hessian (Subsection 3.1). We then expand the kernel using the Raney machinery from Section 2, obtaining an exact Gram representation as a sum of rank-one operators (Subsection 3.3). Finally, we analyze the Gram weights and locate the analytic threshold $\zeta_{c}$ at which the naive unweighted Gram realization breaks down (Subsection 3.4).
@@ 3.1. Kernel representation and scale-invariant formulation
The dispersionless Toda hierarchy associates to a simply connected domain $D\subset\mathbb{C}$ a $\tau$-function whose logarithm $\mathcal{F}=\log\tau$ serves as a generating function for the harmonic moments. In the present paper, however, we use this framework only for the polynomial family (1.3).
For the maps (1.3) with $0<\zeta<\zeta_{\mathrm{univ}}$, the boundary is real-analytic and the Wiegmann–Zabrodin kernel identity applies. The mixed second derivatives of $\mathcal{F}$ with respect to the Toda times $(t_{m},\bar{t}_{n})$ are then expressed via the inverse conformal map $w(z)$ as follows [25, 12]:
@@
The scale-invariant Hessian $H=(H_{mn})_{m,n\geq 1}$ is naturally a positive semidefinite quadratic form on finitely supported sequences. Spectral statements require a compatible Hilbert realization, and the standard $\ell^{2}(\mathbb{N})$ topology is not stable at analytic criticality. The appropriate framework is the weighted Gram realization introduced later in Definition 4.1. All spectral statements in Sections 45 refer to these renormalized operators.
@@ 3.3. Gram representation
Then the scale-invariant Hessian admits the exact entrywise representation
@@
where $\bm{v}\otimes\bm{v}^{*}$ denotes the rank-one operator $(\bm{v}\otimes\bm{v}^{*})_{mn}=v_{m}\overline{v_{n}}$. Equivalently, (3.9) holds as a quadratic-form identity on finitely supported sequences.
@@
The Gram form (3.9) makes positivity of $H$ manifest. Namely, it is a sum of positive semidefinite rank-one operators. More importantly, it separates two sources of complexity: (i) The individual contribution of each logarithmic mode $p$, encoded in the vector $\bm{v}^{(p)}$, and (ii) the collective mixing of modes within each symmetry sector, which determines the eigenvalue distribution.
The spectral behavior of $H$ is controlled by the interplay between these two effects. As we show next, the critical phenomenon at $\zeta=\zeta_{c}$ arises from the borderline summability of the Gram entries, rather than by a singularity in any individual term.
@@ 3.4. Gram weights and the analytic threshold
For each fixed $p\geq 1$, one has $\sigma_{p}(\zeta)<\infty$ if and only if $\zeta<\zeta_{c}$. In particular, the analytic threshold for the Gram weights is exactly $\zeta_{c}$.
@@ 4. Subcritical phase $(0<\zeta<\zeta_{c})$
We now turn to the operator-theoretic picture in the analytic subcritical regime. Proposition 3.5 gives a canonical Gram factorization of each symmetry block of the Hessian. Although the unrenormalized Gram weights $\sigma_{p}(\zeta)$ diverge as $\zeta\uparrow\zeta_{c}$ (Proposition 3.8), an explicit diagonal weighting renormalizes the column modes. The resulting compact positive operators have a single logarithmically diverging eigenvalue, while the remaining soft spectrum stays bounded and admits a well-defined limiting description after compression.
4.1. Operator framework, block factorization, and weighted renormalization
@@
Let $\mathcal{W}:\mathcal{H}_{\beta}\to\ell^{2}(\mathbb{N}_{0})$ be the unitary map:
@@
Identifying $\mathcal{W}$ with the diagonal operator $\operatorname{diag}(w_{j})_{j\geq 0}$ on $\ell^{2}(\mathbb{N}_{0})$ via the canonical basis, the renormalized synthesis operator is
@@
initially defined on finitely supported sequences. Proposition 4.3 shows that it extends uniquely by continuity to a Hilbert–Schmidt operator. Its associated weighted Gram and Hessian operators are
@@
The coefficient matrix of the Hessian block is the intrinsic object. The weights (4.2) neither alter these coefficients nor remove the critical instability. Their role is to place the same coefficients in a fixed $\zeta$-independent Hilbert scale in which the singular part is realized as a compact-plus-rank-one operator.
@@
It is thus enough to bound $\sigma_{p}(\zeta)$ uniformly in $p$. Using the closed form for $R_{s,p}(m)$ from Proposition 2.4, the Gram weights may be written as
@@
for some finite constant $C_{s}^{\prime}(\zeta)$ depending on $\zeta$, but not on $p$. Finally, using $w_{j}^{2}=p_{j}^{3+2\beta}M^{2p_{j}}$ and (4.5),
@@ 4.2. Rank-one logarithmic spike at $\zeta_{c}$
The detailed coefficient analysis is contained in Appendix B. Here we only present the reduction that produces (4.9).
@@
In particular, $R^{(q)}(\zeta)$ is compact for each $\zeta$, and the convergence also holds in operator norm. We denote this operator-norm limit by $\widetilde{C}_{*}^{(q)}$. Thus the $R_{*}^{(q)}$ of Appendix B and the $\widetilde{C}_{*}^{(q)}$ used in the main text are the same operator.
@@
because $R^{(q)}(\zeta)\to\widetilde{C}_{*}^{(q)}$ in Hilbert–Schmidt norm and $L(\zeta)(K_{\eta}-K_{1})\to 0$ in operator norm. Since $K_{1}=\widetilde{\bm{d}}^{(q)}\otimes\widetilde{\bm{d}}^{(q)*}$, this is exactly (4.9). ∎
@@ 4.3. Spectral asymptotics
We now convert the rank-one decomposition of Theorem 4.6 into spectral statement. The conclusion is that the singular part produces one stiff eigenvalue in each symmetry block, while the remaining spectrum stays bounded and converges to the compressed limit operator.
@@
Theorem 4.7 has a clear numerical signature: after weighted renormalization, exactly one eigenvalue in each symmetry block grows on the logarithmic scale $L(\zeta)$, while the remaining eigenvalues stay bounded. In all figures below, $N$ denotes the truncation size of the finite weighted block matrix used for numerical diagonalization. Figure 1 illustrates this separation for the block operator $\widetilde{G}^{(q)}(\zeta)$ in the sector $q=1$ for $s=3$ and $s=5$. The top row shows that the leading eigenvalue $\mu_{1}^{(q)}(\zeta)$ is asymptotically affine in $L(\zeta)$, whereas the bottom row confirms that after division by $L(\zeta)$ only the stiff eigenvalue survives and all soft eigenvalues tend to zero.
Figure 1. Logarithmic spectral asymptotics of the weighted Gram block $\widetilde{G}^{(q)}(\zeta)$ for $s=3,5$ in sector $q=1$ ($\beta=1$, $N=30$). Top row: the leading eigenvalue $\mu_{1}^{(q)}(\zeta)$ plotted against $L(\zeta)$, showing the asymptotically affine law $\mu_{1}^{(q)}(\zeta)=\Gamma^{(q)}\,L(\zeta)+O(1)$. The dashed line is a tail linear fit. Bottom row: the normalized eigenvalues $\mu_{k}^{(q)}(\zeta)/L(\zeta)$. The ratio $\mu_{1}^{(q)}(\zeta)/L(\zeta)$ approaches a positive constant, whereas $\mu_{k}^{(q)}(\zeta)/L(\zeta)\to 0$ for $k\geq 2$.
Proof of Theorem 4.7.
@@
The remaining spectral behaviour is governed by the compression of the limiting remainder to the orthogonal complement of the spike direction.
@@
Proposition 4.8 identifies the limiting bounded spectrum that remains after the stiff direction is removed. This bounded part is most naturally described at the level of the compressed operator, rather than through the full diverging spectrum. Figure 2 illustrates this soft regime for $s=3$ and $s=5$: the upper plots show the convergence of the first soft eigenvalues as $\zeta\uparrow\zeta_{c}$, while the lower plots display near-critical finite-$\zeta$ soft spectral profiles across the symmetry sectors. Together with Figure 1, this makes the stiff/soft decomposition visually explicit.
Figure 2. Soft spectrum after removal of the stiff direction for $s=3,5$ ($\beta=1$, $N=40$). Top row: the soft eigenvalues $\mu_{2}^{(q)}(\zeta),\dots,\mu_{6}^{(q)}(\zeta)$ in sector $q=1$, plotted against $1/L(\zeta)$. Their flattening as $1/L(\zeta)\to 0$ shows that the soft eigenvalues remain bounded as $\zeta\uparrow\zeta_{c}$. Bottom row: finite-$\zeta$ snapshots of the soft eigenvalues for all sectors $q=1,\dots,s$ at $\zeta/\zeta_{c}=0.9999$.
@@
Theorem 4.6, Theorem 4.7, and Proposition 4.8 together yield the content of Theorem A stated in the Introduction.
5. Scalar Gram continuation beyond the analytic threshold